[Brahms-l] Shift report 20040124 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Sat Jan 24 2004 - 07:51:48 EST
------------------------------------------------------------
Date: Sat Jan 24 07:51:47 2004, Shift: 00:00-08:00
Supervisor: ZIS, with: 
Log book#15 pages 41 - 41

Shift Summary:
------------------------------------------------------------
1. Beam from the beggining of the shift until 2:30. 
After that there was three unsuccesfully attempts to 
get beam. The beam stored from 6:00 until the end of shift
2. Run  
   -9837 177K@MRS, 210K@FS
   -9838  94K@MRS, 111K@FS
   -9839 309K@MRS, 366K@FS  
   -9840  (changed to B861) still running at the end of shift
3. No detector problems
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Jan 24 07:52:17 2004

This archive was generated by hypermail 2.1.8 : Sat Jan 24 2004 - 07:52:33 EST